OpenStudy (anonymous):

What is the simplest form of sqrt75x^2/sqrt12xy^2

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

These are some more that I am struggling with

OpenStudy (michele_laino):

starting fro the last: \[\frac{ 7 }{ 2+\sqrt{5} }=\frac{ 7(2-\sqrt{5}) }{ (2+\sqrt{5})(2-\sqrt{5}) }=\] =

OpenStudy (michele_laino):

=\[7\sqrt{5}-14\]

OpenStudy (anonymous):

so would that be f?

OpenStudy (anonymous):

Use the rule that sqrt(xy)=sqrt(x)*sqrt(y) sqrt(75x^5)=sqrt(5)*sqrt(5)*sqrt(3)*sqrt(x^4)*sqrt(x). You can use this to factor out.

OpenStudy (michele_laino):

now: \[(2-\sqrt{7})(1+2\sqrt{7})=2+4\sqrt{7}-\sqrt{7}-14=3\sqrt{7}-12\]

OpenStudy (anonymous):

@ThePackMan that doesn't make sense

OpenStudy (michele_laino):

now: \[2\sqrt{72}-3\sqrt{32}=2\sqrt{9*4*2}-3\sqrt{16*2}=2*3*2\sqrt{2}-3*4\sqrt{2}=0\]

OpenStudy (anonymous):

Ok :)

OpenStudy (michele_laino):

the first one: \[\frac{ \sqrt{75x ^{5}} }{ \sqrt{12xy ^{2}} }=\] \[=\frac{ \sqrt{25*3*x ^{4}*x} }{ \sqrt{4*-3*x*y ^{2}} }=\] \[\frac{ 5x ^{2}\sqrt{x} }{ 2y \sqrt{3x} }=\frac{ 5x ^{2}\sqrt{x} }{ 2\sqrt{3}y \sqrt{x} }=\] \[\frac{ 5x ^{2} }{ 2y \sqrt{3} }\]

OpenStudy (michele_laino):

oops: 75=3*25, so your answer is: \[\frac{ 5x ^{2} }{ 2y }\]

OpenStudy (anonymous):

Perfecto! Have time for 4 more?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Thank you so much! How do I give medals?

OpenStudy (michele_laino):

from n°8: \[g(x)-f(x)=x ^{3}-(-2x+5)=x ^{3}+2x-5\]

OpenStudy (michele_laino):

n°10: \[\frac{ g(x) }{ f(x) }=\frac{ x+3 }{ x ^{2}-2x-15 }=\] =\[=\frac{ x+3 }{ (x-5)(x+3) }\] so domain of g/f= x not equal to 5 and x not equal to -3

OpenStudy (michele_laino):

n°9: \[f*g=3x(x ^{2}+1)=3x ^{3}+3x\]

OpenStudy (michele_laino):

n.7 squaring both sides we have: \[3x-5=16\] so 3x=21, and then x=7 which is acceptable, why?

OpenStudy (anonymous):

X=7 because you divided 21 by 3

OpenStudy (michele_laino):

no, question is more subtle. hint radicand has to be positive, so?

OpenStudy (anonymous):

I have not idea :/